# On the spectrum of Sturmian Hamiltonians of bounded type in a small coupling regime

### v1

@@ 2.1 Torus Model for $\lambda=0$
The surface $S_{0}$ is smooth everywhere except at $P_{1}=(1,1,1),\ P_{2}=(-1,-1,1),\ P_{3}=(1,-1,-1),$ and $P_{4}=(-1,1,-1)$. Let us denote $\mathbb{S}:=S_{0}\cap\{|x_{1}|\leq 1,|x_{2}|\leq 1,|x_{3}|\leq 1\}$. We know
@@
This is a direct calculation. ∎
@@
$\mu^{s}(a):=\frac{a-\sqrt{1+a^{2}}}{2}\ \text{and}\ \mu^{u}(a):=\frac{a+\sqrt{1+a^{2}}}{2}.$
For each $a>1$, the unstable eigenspace $E^{u}_{a}$ lies between the lines $E_{1}^{u}$ and the line $\{y=0\}$.
Let $\beta\in(0.0.1]$, and consider the vector spaces $V_{1}(\beta):=\text{span}\left\{\begin{pmatrix}1\\ -\beta\end{pmatrix}\right\}$ and $V_{2}(\beta):=\text{span}\left\{\begin{pmatrix}1\\ 1+2\beta\end{pmatrix}\right\}$. We let $K^{u}_{\beta}$ be the collection of vectors that are between $V_{1}(\beta)$ and $V_{2}(\beta)$, which indeed contains each $E^{u}_{a}$. Formally,
@@
and we also set $V_{0}(\beta)$ to be the line mid-way between $V_{1}(\beta)$ and $V_{2}(\beta)$. Then, we have the following:
@@
  • $A_{a}^{-1}\left(K^{s}_{\beta}\right)\subset\text{int}K^{s}_{\beta}$ and $A_{a}\left(K^{u}_{\beta}\right)\subset\text{int}K^{u}_{\beta}$
@@
  • $\overline{\mu}_{\beta}|v|\leq\left|A_{a}^{-1}v\right|\leq\left(\mu^{s}(a)\right)^{-1}|v|$ for all $v\in K_{\beta}^{s}$
@@
and hence $\left(a-\beta\right)\begin{pmatrix}1\\ \frac{1}{a-\beta}\end{pmatrix}$ belongs to a line strictly between $V_{1}(\beta)=\text{span}\left\{\begin{pmatrix}1\\ -\beta\end{pmatrix}\right\}$ and $V_{2}(\beta)=\text{span}\left\{\begin{pmatrix}1\\ 1+2\beta\end{pmatrix}\right\}$. A similar calculation shows that the same is true for $A_{a}\begin{pmatrix}1\\ 1+2\beta\end{pmatrix}$. The weakest expansion of any vector from $K^{u}_{\beta}$ under any $A_{a}$ is given by iterating vectors furthest away from $E_{1}^{u}$ under the map $A_{1}$. In other words, for any $a$ and $v\in K^{u}_{\beta}$,
$\left|A_{a}\frac{v}{|v|}\right|\geq\left|A_{1}\frac{v}{|v|}\right|$
$\geq\left|A_{1}\frac{\left(1,-\beta\right)^{T}}{\left|\left(1,-\beta\right)^{T}\right|}\right|=\frac{\sqrt{\left(1-\beta\right)^{2}+1}}{\sqrt{\beta^{2}+1}}=:\overline{\mu}_{\beta}.$
@@
For each $n\in\mathbb{N}$, we consider $\hat{A}_{a_{n}}$ to be the action on $\mathbb{R}\mathbb{P}^{1}$ induced by $A_{a_{n}}$. Denote $\hat{K}^{s}\subset\mathbb{R}\mathbb{P}^{1}$ to be the set obtained from projecting $K^{s}$ onto $\mathbb{R}\mathbb{P}^{1}$, which is compact. Then, from Lemma 2 we have that each map $\hat{A}_{a_{n}}^{-1}$ restricted to $\hat{K}^{s}$ maps $\hat{K}^{s}$ into $\hat{K}^{s}$ and is a contraction with contraction rate no greater than the contraction rate of $\hat{A}_{1}$, which is $\mu^{s}(1)^{2}$. By Proposition 1, there is a $L^{*}\in\hat{K}^{s}$ such that
@@
for any $L\in\hat{K}^{s}$. That is to say, $E^{s}_{\alpha}$ projects uniquely onto $L^{*}$ and is hence a line. In particular, we know that the projection of $\text{span}\left\{(0,1)^{T}\right\}$ onto $\mathbb{R}\mathbb{P}^{1}$, denoted by $\hat{(0,1)}$, lives in $\hat{K}^{s}$, and hence $\lim\limits_{n\rightarrow\infty}\hat{A}_{a_{1}}^{-1}\circ\cdots\circ\hat{A}_{a_{n}}^{-1}\left(\hat{(0,1)}\right)=L^{*}.$ We however see that the line $\text{span}\left\{A_{a_{1}}^{-1}\circ\cdots\circ A_{a_{n}}^{-1}(1,0)^{T}\right\}$ projects onto $\hat{A}_{a_{1}}^{-1}\circ\cdots\circ\hat{A}_{a_{n}}^{-1}\left(\hat{(1,0)}\right)$. That $E^{s}_{\alpha}$ has slope $-\frac{1}{\alpha}$ follows from
@@
Projecting these lines onto $\mathbb{T}^{2}$, we have that each line $\pi\left(E^{s}_{\alpha}\right)$ is dense in $\mathbb{T}^{2}$. If $Q$ is a rational point in $\mathbb{T}^{2}$, and $L\subset\pi\left(E^{s}_{\alpha}\right)$ is a line segment with midpoint 00, the length of $L$ and continued fraction expansion of $\alpha$ determine the distance from $L$ and $Q$. From Proposition 2, we have the following:
@@
Let $C^{*}>0$. By Proposition 2, there is a $n=n\left(C^{*},\mathcal{A}\right)$ so that for any $\alpha=[a_{1},a_{2},\dots]$ where $a_{i}\in\mathcal{A}$, we must have
@@ 2.2 Torus Model for Small $\lambda>0$
    where $F_{\lambda}:=\pi^{-1}_{\lambda}\circ F$.
@@
One can choose $\rho^{\prime}\in(0,\rho)$ such that for $O_{\rho}=\bigcup_{i=1}^{4}B_{\rho}$ and $O_{\rho^{\prime}}=\bigcup_{i=1}^{4}B_{\rho^{\prime}},$ we must have
@@
Denote the bounded component of $S_{\lambda}\setminus O_{\rho}$ by $\mathbb{S}_{\lambda,O_{\rho}}$. Then, there is a $r(\rho)>0$ and an open set $\tilde{O}_{r(\rho)}\subset\mathbb{T}^{2}$ which consists of neighborhoods of $Q_{i}$, $i=1,\dots,4$, each of diameter less than $\frac{r(\rho)}{2}$, such that $F^{-1}\circ\pi_{\lambda}\left(\mathbb{S}_{\lambda,O_{\rho}}\right)=\mathbb{T}^{2}\setminus\tilde{O}_{r(\rho)}$.
Also, there is an open set $O^{\prime}\subset\tilde{O}_{r(\rho)}$ such that $F^{-1}\circ\pi_{\lambda}\left(\mathbb{S}_{\lambda,O_{\rho^{\prime}}}\right)=\mathbb{T}^{2}\setminus O^{\prime}$. Let us denote $F_{\lambda}:=\pi_{\lambda}^{-1}\circ F$.
Let $P\in\mathbb{T}^{2}$. Then, $F_{\lambda}(P)$ is contained in $\mathbb{S}_{\lambda,O_{\rho}}$ and $\left(T_{a}\circ F_{\lambda}\right)(P)\in\mathbb{S}_{\lambda,O_{\rho^{\prime}}}$, so that
@@
Since $\left\|\pi_{\lambda}\circ T_{a}\circ\pi^{-1}_{\lambda}\restriction_{\mathbb{S}_{0,O_{\rho}}}-T_{a}\restriction_{\mathbb{S}_{0,O_{\rho}}}\right\|_{C^{2}}\rightarrow 0$ as $\lambda\rightarrow 0$, it follows that $\left\|\tilde{T}_{\lambda,a}-\hat{T}_{0,a}\right\|_{C^{2}}\rightarrow 0$, on $\mathbb{T}^{2}\setminus O_{r(\rho)}$, as $\lambda\rightarrow 0$.
@@
2.3 Main Problem Setup
@@
2.3.1 Boxes and Local Cones
@@
Remark 2.
Suppose that $a\in\mathbb{N}$ and consider the square $\text{Box}_{1}(0)$. Then, from our choice of $\beta_{0}$ (see 2.2), we know that the length of any segment contained in $K^{u}$ that connects 00 and a boundary line of $\text{Box}_{1}(0)$ parallel to $V^{\perp}$, has length no greater than $\frac{2}{\sqrt{3}}$ and length no less than $1$. If we iterate this segment under $A_{a}$, the its length will be greater than $\overline{\mu}_{\beta_{0}}$. However, from (2.1), we have
@@
This calculation gives that for any line segment $L$ through 00 that is contained in $K^{u}$ and connects the boundary lines of $\text{Box}_{1}(0)$ that are parallel to $V^{\perp}$, we must have that $A_{a}(L)$ is a line through 00 that is contained in $K^{u}$ and extends outside of the square $\text{Box}_{1}(0)$ (see Figure 1).
If $\beta>\beta_{0}$ is close enough to $\beta_{0}$, then same property will hold for $K^{u}_{\beta}$ restricted to $\text{Box}_{1}(0)$. Notice that in this case, $V(\beta_{0})$ is not the center line of the cone $K^{u}_{\beta}$, but the angle between its center line and $V(\beta_{0})$ is small.
The property also holds for a sufficiently small perturbation of each map $A_{a}$.
@@
Figure 1: Illustration of Remark 2
@@
For each $r\in\left(0,\frac{1}{10}\right)$, let us denote $\mathbb{B}_{r}(Q):=\pi\left(\text{Box}_{r}\left(\pi^{-1}(Q)\right)\right)$, which is the rectangular neighborhood centered at $Q$ of side lengths $2r$. We set $\partial^{s}\mathbb{B}_{r}(Q)$ to be the two line segments of $\partial\mathbb{B}_{r}(Q)$ that are parallel to the line $\pi\left(V^{\perp}\right)$ and $\partial^{u}\mathbb{B}_{r}(Q)$ to be the two line segments of $\partial\mathbb{B}_{r}(Q)$ that are parallel to the line $\pi\left(V\right)$ (see Figure 2).
In particular, let us set $\mathbb{B}_{i}:=\mathbb{B}_{\frac{1}{10}}(Q_{i})$, $\partial^{s}\mathbb{B}_{i}:=\partial^{s}\mathbb{B}_{\frac{1}{10}}(Q_{i})$ and $\partial^{u}\mathbb{B}_{i}:=\partial^{u}\mathbb{B}_{\frac{1}{10}}(Q_{i})$.
@@
For any cone $K$ containing $K^{u}$ and transversal to $K^{s}$, we define
@@
For a cone $K\subset\mathbb{R}^{2}$, define the local projected cone $\tilde{K}(Q):=\pi\left(\left(\boldsymbol{Q}+K\right)\cap\text{Box}_{\frac{1}{10}}(\boldsymbol{Q})\right)$. If $K$ is a cone containing $K^{u}$ and transversal to $K^{s}$, then set $\tilde{K}_{+}(Q):=\pi\left(\left(\boldsymbol{Q}+K_{+}\right)\cap\text{Box}_{\frac{1}{10}}(\boldsymbol{Q})\right)$ and $\tilde{K}_{-}(Q):=\pi\left(\left(\boldsymbol{Q}+K_{-}\right)\cap\text{Box}_{\frac{1}{10}}(\boldsymbol{Q})\right)$ (see Figure 3).
@@
With this in mind, we set +$\partial^{s}_{+}\mathbb{B}_{r}(Q)$ to be the line segment from $\partial^{s}\mathbb{B}_{r}(Q)$ that intersects $\tilde{K}_{+}^{u}(Q)$ and $\partial^{s}_{-}\mathbb{B}_{r}(Q)$ to be the line segment from $\partial^{s}\mathbb{B}_{r}(Q)$ that intersects $\tilde{K}_{-}^{u}(Q)$.
We will let $K^{s,u}(x,y)$ be the cone $K^{s,u}$ contained in the tangent space $T_{(x,y)}\mathbb{T}^{2}$, and set
Remark 3.
@@
2.3.2 Dynamical Rectangles

### v2

@@ 2.1 Torus Model for $\lambda=0$
Notice that the surface $S_{0}$ is smooth everywhere except at each $P_{i}$, $i=1,\dots,4$. Let us denote $\mathbb{S}:=S_{0}\cap\{|x_{1}|\leq 1,|x_{2}|\leq 1,|x_{3}|\leq 1\}$. We know
@@
This is a direct calculation which can be found in Lemma 3.15 in [M].
@@
$\mu^{s}(a):=\frac{a-\sqrt{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}4+a^{2}}}}{2}\ \text{and}\ \mu^{u}(a):=\frac{a+\sqrt{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}4+a^{2}}}}{2}.$
For each $a>1$, the unstable eigenspace $E^{u}_{a}$ lies between the lines $E_{1}^{u}$ and the line $\{y=0\}$ in the projective sense.
Let $\beta\in(0{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0},}0.1]$, and consider the vector spaces $V_{1}(\beta):=\text{span}\left\{\begin{pmatrix}1\\ -\beta\end{pmatrix}\right\}$ and $V_{2}(\beta):=\text{span}\left\{\begin{pmatrix}1\\ 1+2\beta\end{pmatrix}\right\}$. We let $K^{u}_{\beta}$ be the collection of vectors that are between $V_{1}(\beta)$ and $V_{2}(\beta)$, which indeed contains each $E^{u}_{a}$. Formally,
@@
and we also set $V_{0}(\beta)$ to be the line mid-way between $V_{1}(\beta)$ and $V_{2}(\beta)$. That is,
$V_{0}(\beta):=\text{span}\left\{\begin{pmatrix}1\\ -\beta\end{pmatrix}+\begin{pmatrix}1\\ 1+2\beta\end{pmatrix}\right\}.$
Then, we have the following:
@@
  • $A_{a}^{-1}\left(K^{s}_{\beta}\right)\subset\text{int}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\left(K^{s}_{\beta}\right)}$ and $A_{a}\left(K^{u}_{\beta}\right)\subset\text{int}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\left(K^{u}_{\beta}\right)}$
@@
  • $\overline{\mu}_{\beta}|v|\leq\left|A_{a}^{-1}v\right|\leq\left({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\left|\mu^{s}(a)\right|}\right)^{-1}|v|$ for all $v\in K_{\beta}^{s}$
@@
and hence $\left(a-\beta\right)\begin{pmatrix}1\\ \frac{1}{a-\beta}\end{pmatrix}$ belongs to a line strictly between $V_{1}(\beta)=\text{span}\left\{\begin{pmatrix}1\\ -\beta\end{pmatrix}\right\}$ and $V_{2}(\beta)=\text{span}\left\{\begin{pmatrix}1\\ 1+2\beta\end{pmatrix}\right\}$ in the projective sense. A similar calculation shows that the same is true for $A_{a}\begin{pmatrix}1\\ 1+2\beta\end{pmatrix}$. The weakest expansion of any vector from $K^{u}_{\beta}$ under any $A_{a}$ is given by iterating vectors furthest away from $E_{1}^{u}$ under the map $A_{1}$. In other words, for any $a$ and $v\in K^{u}_{\beta}$,
$\left|A_{a}\frac{v}{|v|}\right|\geq\left|A_{1}\frac{v}{|v|}\right|\geq\left|A_{1}\frac{\left(1,-\beta\right)^{T}}{\left|\left(1,-\beta\right)^{T}\right|}\right|=\frac{\sqrt{\left(1-\beta\right)^{2}+1}}{\sqrt{\beta^{2}+1}}=:\overline{\mu}_{\beta}.$
@@
Remark 1.
The range of $\beta$ in this lemma can most certainly be increased, but for our arguments, we just need any interval containing 00.
@@
For each $n\in\mathbb{N}$, we consider $\hat{A}_{a_{n}}$ to be the action on real projective space $\mathbb{R}\mathbb{P}^{1}$ induced by $A_{a_{n}}$. Denote $\hat{K}^{s}\subset\mathbb{R}\mathbb{P}^{1}$ to be the set obtained from projecting $K^{s}$ onto $\mathbb{R}\mathbb{P}^{1}$, which is compact. Then, from Lemma 2 we have that each map $\hat{A}_{a_{n}}^{-1}$ restricted to $\hat{K}^{s}$ maps $\hat{K}^{s}$ into $\hat{K}^{s}$ and is a contraction with contraction rate no greater than the contraction rate of $\hat{A}_{1}$, which is $\mu^{s}(1)^{2}$. By Proposition A.1, there is a $L^{*}\in\hat{K}^{s}$ such that
@@
for any $L\in\hat{K}^{s}$. That is to say, $E^{s}_{\alpha}$ projects uniquely onto $L^{*}$ and is hence a line. In particular, we know that the projection of $\text{span}\left\{(0,1)^{T}\right\}$ onto $\mathbb{R}\mathbb{P}^{1}$, denoted by $\hat{(0,1)}$, lives in $\hat{K}^{s}$, and hence $\lim\limits_{n\rightarrow\infty}\hat{A}_{a_{1}}^{-1}\circ\cdots\circ\hat{A}_{a_{n}}^{-1}\left(\hat{(0,1)}\right)=L^{*}.$ We however see that the line $\text{span}\left\{A_{a_{1}}^{-1}\circ\cdots\circ A_{a_{n}}^{-1}(1,0)^{T}\right\}$ projects onto $\hat{A}_{a_{1}}^{-1}\circ\cdots\circ\hat{A}_{a_{n}}^{-1}\left(\hat{(1,0)}\right)$.
@@
This shows that $E^{s}_{\alpha}$ has slope $-\frac{1}{\alpha}$, proving the lemma.
It is clear that each line $\pi\left(E^{s}_{\alpha}\right)$ is dense in $\mathbb{T}^{2}$. If $Q$ is a rational point in $\mathbb{T}^{2}$, and $L\subset\pi\left(E^{s}_{\alpha}\right)$ is a line segment with midpoint 00, the length of $L$ and continued fraction expansion of $\alpha$ determine the distance from $L$ and $Q$. This exact distance can be estimated using Proposition B.1. In particular, we have:
@@
Let $C^{*}>0$. By Proposition B.1, there is a $n=n\left(C^{*},\mathcal{A}\right)$ so that for any $\alpha=[a_{1},a_{2},\dots]$ where $a_{i}\in\mathcal{A}$, we must have
@@ 2.2 Torus Model for Small $\lambda>0$
    where $F_{\lambda}:=\pi^{-1}_{\lambda}\circ F$ and $F$ is defined in Lemma 1.
@@
One can choose $\rho^{\prime}\in(0,\rho)$ such that for $O_{\rho}=\bigcup_{i=1}^{4}B_{\rho}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\left(P_{i}\right)}$ and $O_{\rho^{\prime}}=\bigcup_{i=1}^{4}B_{\rho^{\prime}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\left(P_{i}\right)},$ we must have
@@
Denote the bounded component of $S_{\lambda}\setminus O_{\rho}$ by $\mathbb{S}_{\lambda,O_{\rho}}$. Then, there is a $r(\rho)>0$ and an open set $\tilde{O}_{r(\rho)}\subset\mathbb{T}^{2}$ which consists of neighborhoods of $Q_{i}$, $i=1,\dots,4$, each of diameter less than $\frac{r(\rho)}{2}$, such that $F^{-1}\circ\pi_{\lambda}\left(\mathbb{S}_{\lambda,O_{\rho}}\right)=\mathbb{T}^{2}\setminus\tilde{O}_{r(\rho)}$. Also, there is an open set $O^{\prime}\subset\tilde{O}_{r(\rho)}$ such that $F^{-1}\circ\pi_{\lambda}\left(\mathbb{S}_{\lambda,O_{\rho^{\prime}}}\right)=\mathbb{T}^{2}\setminus O^{\prime}$.
Let us denote $F_{\lambda}:=\pi_{\lambda}^{-1}\circ F$ and let $P\in\mathbb{T}^{2}$. Then, $F_{\lambda}(P)$ is contained in $\mathbb{S}_{\lambda,O_{\rho}}$ and $\left(T_{a}\circ F_{\lambda}\right)(P)\in\mathbb{S}_{\lambda,O_{\rho^{\prime}}}$, so that
@@
Since $\left\|\pi_{\lambda}\circ T_{a}\circ\pi^{-1}_{\lambda}\restriction_{\mathbb{S}_{0,O_{\rho}}}-T_{a}\restriction_{\mathbb{S}_{0,O_{\rho}}}\right\|_{C^{2}}\rightarrow 0$ as $\lambda\rightarrow 0$, it follows that $\left\|\tilde{T}_{\lambda,a}-\hat{T}_{0,a}\right\|_{C^{2}}\rightarrow 0$, on $\mathbb{T}^{2}\setminus{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{O}_{r(\rho)}}$, as $\lambda\rightarrow 0$.
@@
3 Dynamical Tools and Setup
@@
Namely, in Section 4 we will invoke Lemma 5 and translate Theorem 3 into a problem concerning non-stationary dynamics of Anosov maps around small neighborhoods of each $Q_{i}$. Rather than working with arbitrary neighborhoods, we will define a notion of dynamical rectangles, that are analogous to Markov rectangles for hyperbolic maps, that will be useful in examining the dynamics near each $Q_{i}$. This section is dedicated to precisely defining these rectangles, and studying some basic geometric properties that they exhibit.
3.1 Boxes and Local Cones
Before defining the rectangles with dynamical properties, we will first restrict ourselves to small preliminary boxes around each $Q_{i}$. This subsection is dedicated to choosing these boxes, deriving some needed properties for later, and also introducing some notations.
@@ 3.1 Boxes and Local Cones
For these choice of sets, it is important that each map $A_{a}$ overflows the square $\text{Box}_{1}(0)$. We use this fact later in Theorem 5 and Lemma 16. Formally, we have the following:
Lemma 6.
Let $L$ be a line segment through 00 that is contained in $K^{u}$ and connects the boundary lines of $\text{Box}_{1}(0)$ that are parallel to $V^{\perp}$. Then, for any $a\in\mathbb{N}$, we have that $A_{a}(L)$ is a line through 00 that is contained in $K^{u}$ and extends outside of the square $\text{Box}_{1}(0)$ (see Figure 3).
Moreover, for each $\beta>\beta_{0}$ sufficiently close to $\beta_{0}$, there is a $\epsilon=\epsilon(a,\beta)>0$ such that if $f\in\text{Diff}^{1}\left(\mathbb{R}^{2}\right)$ and $\left\|A_{a}-f\right\|<\epsilon,$ then $f(L)$ is a curve that overflows $\text{Box}_{1}(0)$. That is, $f(L)$ is contained in $K^{u}_{\beta}$, and its endpoints belong to the compliment of $\text{Box}_{1}(0)$.
Proof.
Suppose that $a\in\mathbb{N}$ and consider the square $\text{Box}_{1}(0)$. Then, from our choice of $\beta_{0}$ (see 2.2), we know that the length of any segment $L$ contained in $K^{u}$ that connects 00 and a boundary line of $\text{Box}_{1}(0)$ parallel to $V^{\perp}$, has length strictly less than $\frac{2}{\sqrt{3}}$ and length no less than $1$. If we iterate this segment under $A_{a}$, then its length will be greater than $\overline{\mu}_{\beta_{0}}$. However, from (2.1), we have
@@
It follows that $A_{a}(L)$ is a line through 00 that is contained in $K^{u}$ and extends outside of the square $\text{Box}_{1}(0)$.
From the discussion before (2.1), for $\beta>\beta_{0}$ close enough to $\beta$, we can still guarantee that the length of any segment $L$ contained in $K^{u}_{\beta}$ that connects 00 and a boundary line of $\text{Box}_{1}(0)$ parallel to $V^{\perp}$, has length strictly less than $\frac{2}{\sqrt{3}}$. Then, by shrinking $\beta$ is needed, we have that
$\overline{\mu}_{\beta}\geq(\sqrt{2}-0.01)-0.001$
and hence, this expansion is still greater than $\frac{2}{\sqrt{3}}$. If $L$ is such a segment and $f$ is sufficiently $C^{1}$ close to $A_{a}$, then the endpoints of $f(L)$ will be close to those of $A_{a}(L)$, and we can guarantee that $Df\left(K^{u}_{\beta}\right)\subset K^{u}_{\beta}$. This would imply that the tangent line at each point of $f(L)$ is contained in $K^{u}_{\beta}$, and hence $f(L)\subset K^{u}_{\beta}$. ∎
We now define some sets and notations that will be used frequently throughout the paper.
@@
Figure 3: Illustration of Lemma 6
@@
Definition.
For each $r\in\left(0,\frac{1}{10}\right)$, we define (see Figure 4)
  • $\mathbb{B}_{r}(Q):=\pi\left(\text{Box}_{r}\left(\pi^{-1}(Q)\right)\right)$, which is the rectangular neighborhood centered at $Q$ of side lengths $2r$
  • $\partial^{s}\mathbb{B}_{r}(Q)$ to be the two line segments of $\partial\mathbb{B}_{r}(Q)$ that are parallel to the line $\pi\left(V^{\perp}\right)$
  • $\partial^{u}\mathbb{B}_{r}(Q)$ to be the two line segments of $\partial\mathbb{B}_{r}(Q)$ that are parallel to the line $\pi\left(V\right)$
In particular, let we set $\mathbb{B}_{i}:=\mathbb{B}_{\frac{1}{10}}(Q_{i})$, $\partial^{s}\mathbb{B}_{i}:=\partial^{s}\mathbb{B}_{\frac{1}{10}}(Q_{i})$ and $\partial^{u}\mathbb{B}_{i}:=\partial^{u}\mathbb{B}_{\frac{1}{10}}(Q_{i})$.
@@
Definition.
(see Figure 5) For any cone $K$ containing $K^{u}$ and transversal to $K^{s}$, we define
@@
For a cone $K\subset\mathbb{R}^{2}$, define the local projected cone $\tilde{K}(Q)\subset\mathbb{B}_{\frac{1}{10}}(Q)$ via
$\tilde{K}(Q):=\pi\left(\left(\boldsymbol{Q}+K\right)\cap\text{Box}_{\frac{1}{10}}(\boldsymbol{Q})\right)$
If $K$ is a cone containing $K^{u}$ and transversal to $K^{s}$, then set
$\tilde{K}_{+}(Q):=\pi\left(\left(\boldsymbol{Q}+K_{+}\right)\cap\text{Box}_{\frac{1}{10}}(\boldsymbol{Q})\right)\ \text{and}\ \tilde{K}_{-}(Q):=\pi\left(\left(\boldsymbol{Q}+K_{-}\right)\cap\text{Box}_{\frac{1}{10}}(\boldsymbol{Q})\right)$
With this in mind, we define
  • +$\partial^{s}_{+}\mathbb{B}_{r}(Q)$ to be the line segment from $\partial^{s}\mathbb{B}_{r}(Q)$ that intersects $\tilde{K}_{+}^{u}(Q)$
  • $\partial^{s}_{-}\mathbb{B}_{r}(Q)$ to be the line segment from $\partial^{s}\mathbb{B}_{r}(Q)$ that intersects $\tilde{K}_{-}^{u}(Q)$
  • $K^{s,u}(x,y)$ to be the cone $K^{s,u}$ contained in the tangent space $T_{(x,y)}\mathbb{T}^{2}$
Lastly, we set
@@
The next lemma is another property concerning these boxes that will be frequently used later in Section 4.
Lemma 7.
@@
Proof.
This follows from the fact that any curve tangent to $\mathcal{K}^{u}$ contained in $\mathbb{B}_{\frac{1}{10}}(Q)$ is the graph of a Lipschitz map from $\pi(V)\cap\mathbb{B}_{\frac{1}{10}}(Q)$ to $\pi\left(V^{\perp}\right)\cap\mathbb{B}_{\frac{1}{10}}(Q)$ with Lipschitz constant determined by the cone $K^{u}$. ∎
3.2 Dynamical Rectangles
We are now ready to define these so-called dynamical rectangles.